The following sample observations were randomly selected.




X 4 5 3 6 10




Y 4 6 5 7 7




Determine the regression equation. (6 marks)




i. The following hypotheses are given.






1




: 0




: 0




H




H










=









A random sample of 12 paired observations indicated a correlation of 0.32. Can we




conclude that the correlation in the population is equal to zero? Use 0.05 level of




significance. (3 marks

1
Expert's answer
2021-11-22T09:21:01-0500

The following sample observations were randomly selected.

 

 

 

X 4 5 3 6 10

 

 

 

Y 4 6 5 7 7

 

 

 

Determine the regression equation. (6 marks)

 

 

 

i. The following hypotheses are given.

 

 

 

 

 

1

 

 

 

: 0

 

 

 

: 0

 

 

 

H

 

 

 

H

 

 

 

r

 

 

 

r

 

 

 

=

 

 

 

¹

 

 

 

 

 

A random sample of 12 paired observations indicated a correlation of 0.32. Can we

 

 

 

conclude that the correlation in the population is equal to zero? Use 0.05 level of

 

 

 

significance. (3 marks

 


Answer:


 Correlation quantifies the strength of the linear relationship between a pair of variables, whereas regression expresses the relationship in the form of an equation. 

regression equation is used in stats to find out what relationship, if any, exists between sets of data. For example, if you measure a child’s height every year you might find that they grow about 3 inches a year. 

A linear regression line has an equation of the form Y = a + bX, where X is the explanatory variable and Y is the dependent variable. The slope of the line is b, and a is the intercept (the value of y when x = 0).


X

4

5

3

6

10

Y

4

6

5

7

7

X               Y             x2        XY

                  

4                4          16                 16

5                 6          25                30

3                 5          9                   15

6                  7          36                42

10                 7         100              70

(Total)

28                 29       186            173

(Mean)

5.6                5.8         37.2           34.6


Formula:

    b XY X Y k X Xk a Y bX = − • − = − − = − = − = − = −− = Σ ΣΣ Σ Σ ( )/ (



b= 173-(28)(29)/5 divide by 186 - (28)2/5

=173-(812)/5 divide by 186 – (784)/5

=-644 divide by -29.2

=-22.05


a=5.8 – (-22.5)(5.6)=- 730.8


In the case of random sample above:

t = r=   = square root of 12-2 divide 1-(0.32)2

= square root of 10 divide 0.8976

=11.1408 square root

= 3.3379

a=0.05

We conclude the correlation in the population is equal to zero ,because P value is more than zero.